OpenStudy (anonymous):

What is the quotient: (72y3 + 12y2 - 30y) ÷ 6y

OpenStudy (lgbasallote):

\[\LARGE \frac{72y^3 + 12y^2 - 30y}{6y} \implies \frac{72y^3}{6y} + \frac{12y^2}{6y} - \frac{30y}{6y}\] does that help?

OpenStudy (anonymous):

so would 5 be the quotient

OpenStudy (lgbasallote):

no...it is very far from the answer...

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

So let's just talk about this. Why did you say "so would 5 be the quotient"

OpenStudy (anonymous):

because 30/6 is 5

OpenStudy (anonymous):

OK, but there are three terms, all of them contribute to the answer

OpenStudy (anonymous):

Let's just start with this.. Suppose you had -30 y / 6 y

OpenStudy (anonymous):

it would be -5y

OpenStudy (anonymous):

If you have y on the top (in the "numerator") and on the bottom in the denominator, you know you can cancel them right?

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

So -30 y / 6 y equals ?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Yes. And notice the minus sign. You were missing that before. OK? Now we'll do the other two terms.

OpenStudy (anonymous):

i got 12 and 2

OpenStudy (anonymous):

Well, the first term has y^3 up top and and y below. So you're going to have y^2 left in the answer for that part.. Do you see that?

OpenStudy (anonymous):

So maybe you are a little hung up on the x's and y's???

OpenStudy (anonymous):

yes i see that so would it be 12y^2 + 2y - 5

OpenStudy (anonymous):

OK. So you know the rules. Good. So the question is why did you give your very first answer? Did you just guess? Please tell me so I can help you.

OpenStudy (anonymous):

i divided and cut each exponet down 1

OpenStudy (anonymous):

Your very first answer was so would 5 be the quotient

OpenStudy (anonymous):

Never mind. You know the rules Do the division for numbers Cancel exponents to reduce things like y^2/y Do each term one at a time (separated by + and -) You know what to do!

OpenStudy (anonymous):

thanks for the help